• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Ole Hansen

    Log entry time 00:05:44 on March 17, 2012

    Entry number 366910

    keyword=Swing Shift Sumamry

    16:00 Begin shift waiting for beam for an RHRS optics check
    16:18 Choppy beam arrives. Taking a short run with sieve slits in on both
    16:40 Controlled access to remove sieve slits
    17:05 Access done. Asking for beam.
    17:18 More brief running on thin carbon to check RHRS rate.
    17:45 replay shows that RHRS rates OK. Begin BPM & harp calibration (Alex
    & Pengja)
    19:15 Beam goes away after a power glitch. MCC reports tripped magnets.
    20:03 Beam is back. Resuming BPM & harp calibration.
    22:25 Begin BCM calibration
    22:38 BPM & BCM calibrations done. Begin dilution runs, including an
    on-the-fly rate scan. Experts (Ryan, Melissa) running the show.
    24:00 End shift with dilution runs in progress